OpenStudy (anonymous):

LIMIT SIN(X-2)/(X^2-4) AS X GOES TO 2

OpenStudy (anonymous):

apply la hopitals rule

OpenStudy (anonymous):

= lim cos(x-2) / (2x) = coa(0) / (2(2)) =1/4

OpenStudy (anonymous):

^ second line should be cos(0) / (2(2)) a little typo